# Staircase Cascaded Fusion of Lightweight Local Pattern Recognition and Long-Range Dependencies for Structural Crack Segmentation

### v3

@@ IV-B Implementation Details
2) Comparison Methods: In evaluating model performance, we compared the CrackSCF network with 12 classic segmentation methods and the latest SOTA methods on different datasets. These methods include HED [44], UNet++ [45], Deeplabv3+ [46], AttuNet [47], RCF [48], RIND [49], UCTNet [5], SFIAN [6], CT-crackseg [7], DTrCNet [8], Crackmer [9] and Simcrack [10].
TABLE II: Comparison of results between CrackSCF and other methods on the public dataset and the TUT dataset. Best results are bolded and second best results are underlined
@@
3) Evaluation Metric: To quantitatively evaluate the performance of our proposed crack segmentation model, we used six main metrics: Precision (P), Recall (R), F1 Score ($F1=\frac{2PR}{P+R}$), Mean Intersection over Union (mIoU), Optimal Dataset Scale (ODS), and Optimal Image Scale (OIS). Mean Intersection over Union (mIoU) measures the average ratio of the intersection and union of the predicted results and the true results. The calculation formula for mIoU is:
$MIoU=\frac{1}{i+1}\sum_{a=0}^{i}\frac{p_{aa}}{\sum_{b=0}^{i}p_{ab}+\sum_{b=0}^{i}p_{ba}-p_{aa}}$ (16)
where $i$ is the number of classes, $a$ denotes the true value, $b$ denotes the predicted value, and $p_{ab}$ represents the number of pixels predicted as $a$ for $b$. Given that our crack segmentation task involves binary classification, we set $i=1$.
Optimal Dataset Scale (ODS) indicates the optimal segmentation performance metric obtained when a fixed threshold $t$ is selected over the entire dataset. It is defined as:
$ODS=\max_{t}\frac{2\cdot P_{t}\cdot R_{t}}{P_{t}+R_{t}}$ (17)
Optimal Image Scale (OIS) indicates the optimal segmentation performance metric obtained when an optimal threshold $t$ is selected for each image individually. It is defined as:
$OIS=\frac{1}{N}\sum_{a=1}^{N}\max_{t}\frac{2\cdot P_{t,a}\cdot R_{t,a}}{P_{t,a}+R_{t,a}}$ (18)
where $N$ represents the total number of images.
IV-C Comparison with the SOTA methods
Our proposed method, CrackSCF, was compared with 12 other methods across five public datasets and our own TUT dataset. Table II displays the comparison results for the five public datasets, while Fig. 8 illustrates the visual results of all methods on these datasets.
Figure 8: Visualization results of CrackSCF and other methods on public datasets. Key areas are highlighted with red boxes to indicate easily missed crack pixels and yellow boxes to indicate easily misdetected non-crack areas.
1) Results on the CFD dataset: The crack areas in the CFD dataset are very slender, testing the model’s sensitivity to fine-grained crack pixels. As list in Table II, our proposed CrackSCF attained the best results across all metrics for the CFD dataset. Specifically, on ODS, OIS, P, R, F1, and mIoU, it outperformed the second-best methods by 0.48%, 0.34%, 0.14%, 0.35%, 0.31%, and 0.30%, respectively. Notably, SimCrack [10] achieved the second-best performance on all metrics except for P and R, only behind our method. However, as seen in Fig. 8, its actual segmentation performance is not satisfactory, while our method achieves high-quality segmentation results on various fine-grained crack pixels. Other methods, such as UCTNet [5] and DTrCNet [8], tend to misdetect background noise or miss key detail areas of crack pixels.
2) Results on the GAPS509 dataset: Unlike the CFD dataset, the GAPS509 dataset contains not only finer cracks but also some irrelevant background interference, testing the model’s ability to extract crack areas under high interference conditions. As list in Table II, our method achieved the highest performance on all metrics except R. Although SimCrack [10] achieved 0.6997 on the R metric, leading our method by about 2.72%, its P metric was 0.6171, far lower than our method. The highest F1 score obtained by our method indicates a good balance between P and R. Specifically, our method outperformed the second-best methods on ODS, OIS, P, F1, and mIoU by 1.31%, 1.25%, 2.90%, 1.11%, and 0.45%, respectively. Additionally, as seen in Fig. 8, our method almost perfectly avoids irrelevant background noise, while most other methods falsely detect irrelevant backgrounds and perform worse in segmenting crack pixel areas compared to our method.
3) Results on the DeepCrack dataset: The cracks in the DeepCrack dataset are often relatively wide, testing the model’s ability to extract a large range of crack pixels. As list in Table II, our method achieved the highest results on all metrics for the DeepCrack dataset. Specifically, our method outperformed the second-best methods on ODS, OIS, P, R, F1, and mIoU by 1.08%, 0.39%, 1.51%, 1.04%, 1.42%, and 0.85%, respectively. Notably, CT-crackseg also achieved good results on this dataset, but as seen in Fig. 8, its segmentation performance on detailed cracks is not as good as our method, indicating that our method can effectively handle crack details.
4) Results on the Crack500 dataset: Similar to the DeepCrack, the Crack500 dataset also has relatively wide crack areas, but its structure is complex, posing higher requirements on model performance. As list in Table II, our method achieved the best results on five metrics for the Crack500 dataset, except for R. It outperformed the second-best methods on ODS, OIS, P, F1, and mIoU by 1.11%, 0.07%, 1.69%, 0.47%, and 0.64%, respectively. Although DTrCNet [8] leads on the R metric, our method surpasses it on the P metric by 10.51%. The highest F1 score achieved by our method indicates a good balance between P and R. Additionally, as seen in Fig. 8, our method does not produce false detections in irrelevant background areas, whereas other methods either produce false detections or lack detailed segmentation in crack areas.
5) Results on the CrackMap dataset: Similar to the CFD dataset, the cracks in the CrackMap dataset are also relatively slender. As list in Table II, our method leads all other methods on all metrics, outperforming the second-best methods on ODS, OIS, P, R, F1, and mIoU by 0.48%, 0.34%, 0.14%, 0.35%, 0.31%, and 0.30%, respectively. Notably, SimCrack [10] achieved the second-best results on this dataset, but as seen in Fig. 8, SimCrack’s segmentation performance on the crack tips is not satisfactory, whereas our method effectively identifies these pixels.
@@ IV-C Comparison with the SOTA methods
Figure 9: Visualization results of CrackSCF and other methods on TUT dataset. Key areas are highlighted with red boxes to indicate easily missed crack pixels and yellow boxes to indicate easily misdetected non-crack areas.
TABLE III: Comparison of the complexity of CrackSCF with other methods. Best results are bolded and second best results are underlined
@@
6) Results on the TUT dataset: Fig. 9 shows the visual results of all methods on the TUT dataset. We manually highlighted easily missed crack pixel areas with red boxes and misdetected non-crack pixel areas with yellow boxes. As list in Table II, our method achieved excellent results on the TUT dataset, leading in all metrics. Specifically, it outperformed the second-best methods on ODS, OIS, P, R, F1, and mIoU by 2.73%, 2.18%, 0.98%, 0.51%, 2.12%, and 1.67%, respectively. SimCrack [10] achieved the second-best results on ODS, OIS, F1, and mIoU, but as seen in Fig. 9, its segmentation performance on some key details has room for improvement compared to our method.
TABLE IV: Comparison of CrackSCF using LRDS blocks with networks using other convolutions. Best results are bolded and second best results are underlined
@@
TABLE V: Comparison of CrackSCF Networks using SCFM, signle Concat branch and signle PAF branch. Best results are bolded and second best results are underlined
Method ODS OIS P R F1 mIoU FLOPs ↓ Params ↓ FPS ↑
Signle Concat branch 0.8019 0.8092 0.8228 0.8209 0.8219 0.8359 7.21G 4.71M 42
Signle PAF branch 0.8137 0.8202 0.8282 0.8297 0.8293 0.8427 8.24G 4.72M 39
SCFM 0.8202 0.8266 0.8289 0.8484 0.8382 0.8473 9.26G 4.79M 36
TABLE VI: Comparison of SCFM performance with other segmentation heads. Best results are bolded and second best results are underlined
@@
TABLE VII: Comparison of the performance of using our lightweight deformable attention in LDE with other variants. Best results are bolded and second best results are underlined
@@ IV-D Complexity Analysis
Table III presents our method’s performance alongside other comparisons in terms of FLOPs, Params, and FPS. Our approach achieves notably low values in FLOPs (9.26G) and Params (4.79M), surpassing the second-best by 38.01% and 18.81%, respectively. This underscores our algorithm’s efficiency, making it suitable for implementation on resource-constrained edge devices. In FPS, our model achieves the second-best performance, processing 36 frames per second, trailing only behind the speed-focused DTrCNet [8]. This highlights our model’s robust real-time capability, crucial for prompt crack image segmentation in practical applications. Overall, our model combines low parameters and computational load with competitive processing speed, ensuring effective deployment on devices like smartphones, drones, and other portable platforms.
IV-E Ablation Studies
To evaluate the strengths of our proposed module over other modules, we performed ablation experiments on the TUT dataset.
@@ IV-E Ablation Studies
1) Ablation study of different convolution types: As list in Table IV, on six metrics, our method outperforms methods that use depth-separable convolution or simple low-rank approximate convolution. Our method reduces FLOPs by 61.26% and 56.69%, respectively, and Params by 0.16M compared to the simple low-rank approximation convolution, achieving the highest FPS. Compared to original convolution operations, our method reduces FLOPs and Params by 89.70% and 73.19%, respectively, without degrading performance. These results indicate that LRDS effectively reduces computational load and utilizes parameters while maintaining or improving segmentation performance, enhancing processing speed and benefiting lightweight model design.
2) Ablation study in SCFM using different branch: To demonstrate the effectiveness of SCFM in fusing local detail features and long-range information, we performed ablation experiments on the TUT dataset, as list in Table V. Our method, using Concat and PAF dual-branch feature fusion, achieves the best results across all six segmentation metrics. Using only the PAF branch results in a 1.06% drop in the F1 score. While SCFM slightly increases parameter count, computational load, and inference speed, it significantly improves segmentation performance. These results indicate that our module effectively balances performance and efficiency.
@@
3) Ablation study of different segmentation heads: To demonstrate the advantages of SCFM over other high-performance segmentation heads, we conducted ablation experiments on different segmentation heads, including PID [50], SegFormer [51], UNet [52], and Hamburger head [53], using the TUT dataset. As list in the Table VI, SCFM achieves the best results in segmentation performance metrics. Particularly in terms of F1 and mIoU, SCFM outperforms the UNet head, which achieved the second-best results, by 4.35% and 3.05%, respectively. This indicates that SCFM effectively enhances the model’s segmentation performance by perceiving and integrating local details and long-range dependencies of crack pixels. Furthermore, SCFM requires low computational resources, with Params and FLOPs being 11.95% and 68.84% lower than those of the second-best results, respectively, and achieves the highest FPS. This demonstrates that our SCFM maintains excellent performance while requiring low computational resources.
@@
4) Ablation study of different attention variants in LDE: To validate that our LR Linear can effectively reduce the computational resources required by the deformable attention (MSDA) in LDE while achieving excellent segmentation performance, we compared lightweight MSDA with common multi-head attention (MHA), MHA with embedded LR Linear, and common MSDA. As shown in the Table VII, our method achieves the best results in segmentation performance metrics. Specifically, compared to the common MSDA, our method improves by 1.51% and 1.19% in F1 and mIoU, respectively. Compared to MHA, MSDA outperforms it in segmentation performance, demonstrating the effectiveness of our use of MSDA. Furthermore, whether for MHA or MSDA, when embedded with LR Linear, the required Params and FLOPs significantly decrease. For MSDA, the Params and FLOPs are reduced by 43.43% and 17.70%, respectively. This demonstrates that our LR Linear can effectively reduce the computational requirements needed by the attention mechanism.
V Conclusion

### v4

@@ 4.2 Implementation Details
Table 2: Comparison of results between CrackSCF and other methods on the public dataset and the TUT dataset. Best results are bolded and second best results are underlined
@@
2) Comparison Methods: To evaluate the model performance, we compared the CrackSCF with seven classical and latest SOTA methods on different datasets. These methods include RIND [26], UCTNet [8], SFIAN [11], CTCS [9], DTrCNet [7], Crackmer [10] and Simcrack [27].
@@
3) Evaluation Metric: To quantitatively evaluate the performance of our proposed crack segmentation model, we used six main metrics [11]: Precision (P), Recall (R), F1 Score ($F1=\frac{2PR}{P+R}$), mean Intersection over Union (mIoU), Optimal Dataset Scale (ODS), and Optimal Image Scale (OIS).
4.3 Comparison with the SOTA methods
1) Results on the our TUT dataset: Fig. 6 shows the visual results of all methods on the TUT dataset. We manually highlighted easily missed crack pixel areas with red boxes and misdetected non-crack pixel areas with yellow boxes. As list in Table 2, our method achieved excellent results on the TUT dataset, leading in all metrics. Specifically, it outperformed the second-best methods on ODS, OIS, P, R, F1, and mIoU by 2.73%, 2.18%, 0.98%, 0.51%, 2.12%, and 1.67%, respectively. SimCrack [27] achieved the second-best results on ODS, OIS, F1, and mIoU, but as seen in Fig. 6, its segmentation performance on some key details has room for improvement compared to our method.
2) Results on other public datasets: The CFD, CrackMap and GAPS509 datasets are characterized by slender and fine-grained crack structures, testing the model’s sensitivity to subtle and narrow crack pixels. As listed in Table 2, our proposed CrackSCF achieved the best performance on the CFD and CrackMap datasets. On both datasets, CrackSCF outperforms Simcrack by an average of 1.97%, 9.89%, 2.85%, 1.83%, and 1.00% on ODS, OIS, P, F1, and mIoU respectively. Although SimCrack [27] ranked second on most metrics, its actual segmentation performance on fine crack pixels and crack tips was unsatisfactory, as shown in Fig. 6. Other methods such as UCTNet [8] and DTrCNet [7] often failed to distinguish crack regions from background noise or missed crucial fine-grained crack details.
On the more challenging GAPS509 dataset, our method also achieved the highest performance on all metrics except recall. SimCrack obtained a higher recall value of 0.6997, about 2.72% higher than our method, but its precision was significantly lower at 0.6171, leading to a lower F1 score overall. Our method maintained a superior balance between precision and recall, with improvements of 1.31%, 1.25%, 2.90%, 1.11%, and 0.45% on ODS, OIS, P, F1, and mIoU, respectively. Visual results in Fig. 6 further confirm that our method effectively suppresses false detections in irrelevant background areas while preserving accurate segmentation of slender cracks.
@@ 4.3 Comparison with the SOTA methods
Figure 6: Visualization results of CrackSCF and other methods on TUT and other public dataset. Key areas are highlighted with red boxes to indicate easily missed crack pixels and yellow boxes to indicate easily misdetected non-crack areas.
Both DeepCrack and Crack500 datasets are characterized by relatively wide crack regions, testing the model’s capability to capture extensive crack areas. As listed in Table 2, our method achieved the best results across all metrics on the DeepCrack dataset, outperforming the second-best methods by 1.08%, 0.39%, 1.51%, 1.04%, 1.42%, and 0.85% in terms of ODS, OIS, P, R, F1, and mIoU. Although CTCS [9] also performed well, it fell short in preserving crack detail segmentation, as illustrated in Fig. 6. This demonstrates the advantage of our method in accurately segmenting both wide crack regions and their finer structures. On the Crack500 dataset, our method achieved leading performance in five of the six evaluation metrics: ODS, OIS, P, F1, and mIoU, surpassing the second-best methods by 1.11%, 0.07%, 1.69%, 0.47%, and 0.64%. While DTrCNet [7] achieved a higher recall, our method outperformed it in precision by a notable 10.51%, leading to the highest F1 score and indicating a superior balance between detecting true crack pixels and suppressing false positives. Furthermore, as shown in Fig. 6, our method produces fewer erroneous detections in irrelevant background areas and achieves more accurate and refined segmentation of complex crack structures compared to competing approaches.
Table 3: Comparison of the complexity of CrackSCF with other methods
@@
Table 4: Comparison of CrackSCF using LRDS blocks with networks using other convolutions
@@
Table 5: Comparison of SCFM performance with other segmentation heads
@@
Table 6: Comparison of the performance of using our lightweight deformable attention in LDE with other variants
@@ 4.4 Complexity Analysis
Table 3 presents our method’s performance alongside other comparisons in terms of FLOPs, Params, and FPS. Our approach achieves notably low values in FLOPs at 9.26G and Params at 4.79M, surpassing the second-best by 38.01% and 18.81%, respectively. This underscores our algorithm’s efficiency, making it suitable for implementation on resource-constrained edge devices. In FPS, our model achieves the second-best performance, processing 36 frames per second, trailing only behind the speed-focused DTrCNet [7]. This highlights our model’s robust real-time capability, crucial for prompt crack image segmentation in practical applications. Overall, our model combines low parameters and computational load with competitive processing speed, ensuring effective deployment on devices like smartphones, drones, and other portable platforms.
4.5 Ablation Studies
We conducted ablation studies on the TUT dataset to assess the effectiveness of our proposed module.
@@ 4.5 Ablation Studies
1) Ablation study of different convolution types: As list in Table 4, on six metrics, our method outperforms methods that use depth-separable convolution or simple low-rank approximate convolution. Our method reduces FLOPs by 61.26% and 56.69%, respectively, and Params by 0.16M compared to the simple low-rank approximation convolution, achieving the highest FPS. Compared to original convolution operations, our method reduces FLOPs and Params by 89.70% and 73.19%, respectively, without degrading performance. These results indicate that LRDS effectively reduces computational load and utilizes parameters while maintaining or improving segmentation performance, enhancing processing speed and benefiting lightweight model design.
@@
2) Ablation study of different segmentation heads: To demonstrate the advantages of SCFM over other high-performance segmentation heads, we conducted ablation experiments on different segmentation heads, including PID [28], SegFormer [6], UNet [29], and Hamburger head [30], using the TUT dataset. As list in the Table 5, SCFM achieves the best results in segmentation performance metrics. Particularly in terms of F1 and mIoU, SCFM outperforms the UNet head, which achieved the second-best results, by 4.35% and 3.05%, respectively. This indicates that SCFM effectively enhances the model’s segmentation performance by perceiving and integrating local details and long-range dependencies of crack pixels. Furthermore, SCFM requires low computational resources, with Params and FLOPs being 11.95% and 68.84% lower than those of the second-best results, respectively, and achieves the highest FPS. This demonstrates that our SCFM maintains excellent performance while requiring low computational resources.
@@
3) Ablation study of different attention variants in LDE: To validate that our LR Linear can effectively reduce the computational resources required by the deformable attention (MSDA) in LDE while achieving excellent segmentation performance, we compared lightweight MSDA with common multi-head attention (MHA), MHA with embedded LR Linear, and common MSDA. As shown in the Table 6, our method achieves the best results in segmentation performance metrics. Specifically, compared to the common MSDA, our method improves by 1.51% and 1.19% in F1 and mIoU, respectively. Compared to MHA, MSDA outperforms it in segmentation performance, demonstrating the effectiveness of our use of MSDA. Furthermore, whether for MHA or MSDA, when embedded with LR Linear, the required Params and FLOPs significantly decrease. For MSDA, the Params and FLOPs are reduced by 43.43% and 17.70%, respectively. This demonstrates that our LR Linear can effectively reduce the computational requirements needed by the attention mechanism.
5 Conclusion